OpenStudy (anonymous):

Simplify 4*sq rt of 8 x 5*sq rt of 3 a. 240* sq rt of 2 b. 80* sq rt of 6 c. 120* sq rt of 2 d. 40* sq rt of 6

OpenStudy (anonymous):

B ?

OpenStudy (unklerhaukus):

\[4\sqrt 8 \times 5\sqrt3\]\[=(4\times5)\times\sqrt{8\times3}\]\[=\cdots\]

OpenStudy (anonymous):

97.97958971132712

OpenStudy (unklerhaukus):

dont approximate the irrational component

OpenStudy (anonymous):

Don't calculate any decimals, just multiply the 4x5 then the 8x3. That gives you \[20\sqrt{24}\], right? How could you simplify the square root of 24 even more to get one of the answers you listed? There are several ways to split it up -- \[\sqrt{24}=\sqrt{2}*\sqrt{12}=\sqrt{3}*\sqrt{8}=\sqrt{4}*\sqrt{6}\] , etc. Which way of splitting it up is the best that can be simplified to give you one of the answers you can choose from?

OpenStudy (anonymous):

\[\huge \color{green} {4\sqrt{8} \times 5\sqrt{3} = 20\sqrt{8 \times 3} = 20 \times 2\sqrt{2 \times 3}}\]

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

What don't you understand, how to split up the square root of 24 into two square roots like 2 and 12, or 3 and 8, or 4 and 6?

OpenStudy (anonymous):

Or do you not understand what to do after that?

OpenStudy (anonymous):

all of it. i end up getting weird answers when i try solving them

OpenStudy (anonymous):

Well, show your steps so we can see where you get hung up. Starting from the beginning, how do you start to simplify?

OpenStudy (anonymous):

UnkleRhaukus already did step one for you. What next?

OpenStudy (anonymous):

i aldrady gave him the answer i came up with

OpenStudy (anonymous):

Don't calculate out any decimals, just keep simplifying. The answers you can choose from all don't involve any decimals, so don't do that. Look at what I said before, about how you can split the square root of 24 up. 24 comes from doing 1 * 24, or 2 *12, or 3*8, or 4*6. Let's try splitting up the 24 into the 4 and 6: \[20\sqrt{24}=20\sqrt{4*6}=20\sqrt{4}\sqrt{6}\] Okay so far?

OpenStudy (zzr0ck3r):

we cant tell whats wrong from just an answer, we need to make sure you did it in a logical way

OpenStudy (anonymous):

so i solve this problem by doing the square root of 4 times t he square root of 6 times 20 ? @mbradar2

OpenStudy (eyust707):

correct

OpenStudy (anonymous):

and i got the same answer as before 97.97958971132712. how does that help me solve this ? ughhh !

OpenStudy (anonymous):

Again, don't calculate out any decimals, your final answer will have square roots. So we have \[20*\sqrt{4}*\sqrt{6}\] so far, right? What is the square root of 4 equal to?

OpenStudy (anonymous):

The answer is just a step or two away.